OpenStudy (anonymous):

how do you find the intercepts for x-1=y

jigglypuff314 (jigglypuff314):

the y-intercept is when x = 0 the x-intercept is when y = 0 so to find the y-intercept, set x to 0 so for x-1=y => (0) - 1 = y simplify and that's your y-intercept to find the x-intercept, set y to 0 so for x-1=y => x - 1 = (0) solve for x +1 +1 x = 1

OpenStudy (anonymous):

^